Please help me solve this problem, i have to graph it.
25x^2+4y^2=100

Answer by scott8148(5896)

the second order x and y terms have coefficients that are different, and all the terms are positive
___ this is an ellipse

since there are no first order terms, the ellipse is centered at the origin (0,0)

when x=0 ___ 4y^2=100 ___ y^2=25 ___ y=±5 (these are the y-intercepts)

when y=0 ___ 25x^2=100 ___ x^2=4 ___ x=±2 (these are the x-intercepts)